2 times the square root of x+2 - the square root of x-3 = the square root of x+9
I'm really confused. Would you help me please?

Answer by RAY100(1637)

2(sq rt (x+2)) - sqrt(x-3) = sqrt(x+9)
square both sides
4(x+2) -4 sq rt(x-3) * sqrt (x+2) + (x-3) = (x+9)
move all to rt side except sq rt 's
-4 sq rt (x-3) * sqrt (x+2) = (x+9) -4(x+2) -(x-3)= x+9-4x-8-x+3= -4x+4
divide by (-4) and combine sq rt's
sqrt (x-3) * (x+2) = (x-1)
now square both sides again
(x-3) * (x+2) = (x^2 -2x +1)
x^2 -x -6 = x^2 -2x +1
add x+6 to both sides
0= -x +7
x+7 ANSWER
checking initial expression
2sqrt9 -sqrt4=sqrt16
6-2=4
4=4 ok (ignoring sqrt = (- numbers)